A container is filled with an ideal gas to a pressure of 11.0 atm at 0oC.

  a. What will be the pressure in the container if it is heated to 45o C?


1
Expert's answer
2021-04-16T04:58:35-0400

From the ideal gas equation;

"P1V1\over T1" "=" "P2V2\over T2"

"P1=11.0atm"

"T1=(0+273K)"

"V1=22.4L(STP )"

"P2=unknown"

"T2=(273+45)"

"V2=22.4L(STP)"

Applying the formula;

"P2=" "P1V1T2\over V2T1"

"P2=" "11.0\times \cancel {22.4}\times 318\over \cancel{22.4}\times 273" "=12.81atm"

"\therefore P2=12.81atm"